# Vibration-based damage detection of a trainer jet via multiple input tangential interpolation

### v3

@@ 2. Mode #2 2nd bending (downwards);
MAC Values [-] – main diagonal terms
Case Mode 1 Mode 2 Mode 3
2 0.99 1.00 0.99
3 0.92 1.00 0.97
4 0.86 0.99 0.98
5 0.72 1.00 0.93
6 0.90 0.96 0.99
7 0.85 1.00 0.98
8 0.58 0.99 0.98
9 0.80 0.99 0.99
10 0.88 0.91 0.99
11 0.60 0.91 0.99
12 0.85 0.90 0.99
@@ 4.2.1 Experimentally simulated damage
In this subsubsection, the simulated damage conditions (via mass addition), corresponding to cases #2-7, are investigated. In Figure 15, the MTMAC values for these cases are shown. A relationship exists between the different mass quantities and positions. In fact, cases #4 and 5 show the highest MTMAC values of the six cases. Hence, there is a more pronounced modal parameters change with the baseline case. Traditionally, this is interpreted as a location within the structure that is more sensitive to damage effects.
@@
Now, by looking at all cases with the same mass (M3 in cases #3, 5, and 7), it can be asserted that the location PW_RLE is the most sensitive to damage and that PW_CTE is the least, while PW_TLE lays in the middle. Nevertheless, a deviation from the baseline $\omega_{1-3}$ and $\bm{\phi}_{1-3}$ can be identified in all three locations; most importantly, the magnitude of the added mass influences them by increasing the deviations.
@@
Figure 16: BAE Systems Hawk T1A aircraft: Scaled COMAC for the port wing identified via iLF from cases #1 (Figure 16(a)), 2 (Figure 16(b)), 3 (Figure 16(c)), 4 (Figure 16(d)), 5 (Figure 16(e)), and 6 (Figure 16(f)).
Cases #2 and 3 (Figures 16(a) and 16(b)): from the two images, it is clear that the areas most influenced by the added mass near the centre of the trailing edge (TE) are near the tip and closer to the root at $z\approx$2 m. Notably, as the mass is increased from M1 to M3, the deviation in the mode shapes becomes more dominant near the tip, while retaining some deviation at the same near root area.
@@
Cases #6 and 7 (Figures 16(e) and 16(f)): the most prominent areas where a deviation in $\bm{\phi}_{1-3}$ are seen are the near-tip region and the same close to the root section in cases #2 and 3. Notably, the maximum deviation for both cases is seen near the damaged area. However, when the mass is increased from M2 to M3, the maximum deviation point shifts slightly towards the root (from $x\approx$4 m to -3.5 m).
@@ 4.2.2 Real, panel-off, damage
In Figure 17, the MTMAC values for the panel-off cases (#8-12) are shown. Notably, the values for the panels removed at PW2-3 and 5 have a similar magnitude, while for the PW1 and PW4 cases, the MTMAC values are doubled. This larger deviations in $\omega_{1-3}$ and $\bm{\phi}_{1-3}$ from case #1 show that the latter locations have a higher sensitivity to damage.
@@
Figure 18 shows the scaled COMAC for all of the panel-off cases. It is important to see that for all different locations, different deviations are identified with the COMAC; however, the maximum COMAC never exactly matches the damaged area. This will be discussed later in the paper.
@@
Figure 18: BAE Systems Hawk T1A aircraft: Scaled COMAC for the port wing identified via iLF from cases #10 (Figure 18(a)), 11 (Figure 18(b)), 12 (Figure 18(c)), 13 (Figure 18(d)), and 14 (Figure 18(e)).
5 Discussion
@@
For all cases, added mass and panel-off, the MTMAC successfully detects a change in the system modal parameters $\omega_{1-3}$ and $\bm{\phi}_{1-3}$. In particular, from the two subsets (added mass and panel-off), it is possible to see what locations are more sensitive to changes in the structure. The highest MTMAC values are recorded for cases #8 and 11, which, respectively, represent panel-off cases at locations near the root (PW1) and at mid-span (PW4). The other three panel-off cases show a similar MTMAC value, hence showing a similar sensitivity. Notably, all panels are roughly placed between the quarter chord and mid-chord. On the other hand, the added mass locations are close to either the LE or TE. For them, the most sensitive location to damage is case #5, near the root leading edge. In contrast, the location least sensitive to damage is at TE at mid-span. Nevertheless, in all cases, it is possible to identify the increase in the added mass with an increase in the MTMAC, as shown in Figure 15.
@@
With respect to damage localisation, this is directly achieved only for cases #5-7. Nevertheless, each couple of added mass cases is somewhat coherent in their COMAC values distribution and, most importantly, it is unequivocally different from the others. The latter also holds for the panel-off cases. These findings indicate that the COMAC can be effectively utilised for damage localisation in complex structures, particularly when paired with a high-fidelity finite element model of the system under examination or an empirical measure to correlate the COMAC value distribution with the predicted damage location. Unfortunately, neither was available for the wing under scrutiny at the time of writing; this aspect is not addressed in this work, as it lies beyond its scope, i.e., introducing the use of iLF for MTMAC-based SHM in MIMO tests.
@@ 6 Conclusions
In this work, a practical application of control, in the form of system identification for model parameter estimation, is proposed for damage detection and localisation on a complex aeronautical system. A two-step procedure, including modal parameters identification with the improved Loewner Framework (iLF) and damage detection and quantification with the modified total modal assurance criterion (MTMAC), has been introduced. This is the first instance that the iLF has been used within structural health monitoring and that the MTMAC has been employed as a standalone damage assessment index. Furthermore, as a first, damage detection and localisation have been carried out on the multi-input multi-output vibration dataset of the BAE Systems Hawk T1A aircraft from the Laboratory for Verification and Validation at the University of Sheffield. These resulted in the following findings:
@@
    The iLF has been numerically and experimentally proven to detect small damage-related changes in modal parameters from multi-input multi-output (MIMO) case studies - either numerically-simulated multiple degrees of freedom systems and complex experimental data (the Hawk T1A trainer jet);
@@
    Damage was successfully identified in 11 cases of the recently introduced Hawk T1A aircraft MIMO dataset, a first in the literature.
@@ Appendix A Euler-Bernoulli beam mass and stiffness matrices
Conceptualisation, G.D., and M.C.; methodology, G.D., and M.C.; software, G.D.; validation, G.D. and M.C.; formal analysis, G.D.; investigation, G.D. and M.C.; resources, G.D, M.C., A.M., B.C.; data curation, G.D. and M.C..; writing—original draft preparation, G.D. and M.C.; writing—review and editing, G.D, M.C., A.M., B.C., and OE.BM.; visualisation, G.D., M.C, and A.M.; funding acquisition, G.D, M.C., B.C., A.M., and OE.BM..
@@
The first author is supported by the Grants for research activity of young PhD holders, part of the Universidad Carlos III de Madrid (UC3M) Own Research Program (Ayudas para la Actividad Investigadora de los Jóvenes Doctores, del Programa Propio de Investigación de la UC3M). Furthermore, the authors from UC3M disclosed receipt of the following financial support for the research, authorship, and/or publication of this article: This work has been supported by the Madrid Government (Comunidad de Madrid – Spain) under the Multiannual Agreement with the UC3M (IA_aCTRl-CM-UC3M). The second and fourth authors are supported by the Centro Nazionale per la Mobilità Sostenibile (MOST – Sustainable Mobility Center), Spoke 7 (Cooperative Connected and Automated Mobility and Smart Infrastructures), Work Package 4 (Resilience of Networks, Structural Health Monitoring and Asset Management). The third author gladly acknowledges funding as Beatriz Galindo Distinguished Senior Researcher by the Spanish Government.
@@ Acknowledgements
The experimental and numerical results (Data file 1) supporting this work are openly available from the Zenodo repository at [https://zenodo.org/records/13866467]. In addition, this study used third-party experimental data (Data file 2) made available at [orda.shef.ac.uk/articles/dataset/BAE_T1A_Hawk_Full_Structure_Modal_Test/24948549] under licence that the authors do not have permission to share:
@@
    Data file 1. Data supporting: Vibration-based damage detection of a trainer jet via multiple input tangential interpolation;
@@
    Data file 2. BAE T1A Hawk Full Structure Modal Test

### v4

@@ 2. Mode #2: 2nd bending (downwards);
Mode # Case 2 3 4 5 6 7 8 9 10 11 12
1 0.99 0.92 0.86 0.72 0.90 0.85 0.58 0.80 0.88 0.60 0.85
2 1.00 1.00 0.99 1.00 0.96 1.00 0.99 0.99 0.91 0.91 0.90
3 0.99 0.97 0.98 0.93 0.99 0.98 0.98 0.99 0.99 0.99 0.99
@@ 4.2.1 Experimentally simulated damage
In this subsubsection, the simulated damage conditions (via mass addition), corresponding to cases #2-7, are investigated. In Figure 15, the MTMAC values for these cases are shown. A relationship exists between the different mass quantities and positions. In fact, cases #4 and 5 show the highest MTMAC values of the six cases. Hence, there is a more pronounced change in modal parameters with the baseline case. Traditionally, this is interpreted as a location within the structure that is more sensitive to damage effects.
@@
Now, by looking at all cases with the same mass (M3 in cases #3, 5, and 7), it can be asserted that the location PW_RLE is the most sensitive to damage and that PW_CTE is the least, while PW_TLE lies in the middle. Nevertheless, a deviation from the baseline $\omega_{1-3}$ and $\bm{\phi}_{1-3}$ can be identified in all three locations; most importantly, the magnitude of the added mass influences them by increasing the deviations.
In this sense, it is possible to compare these findings to the threshold defined before for the numerical case study. As mentioned there, establishing reliable thresholds would require a larger dataset of repeated measurements under varying environmental/operational conditions, and well-defined requirements regarding acceptable margins of stiffness or frequency reduction. However, just to showcase the generalisability of the findings across datasets, one can notice that:
  • 1.
    across all locations, smaller added masses (M1, $\approx 250$ g) produce MTMAC values in the same order of magnitude ($\approx 0.04–0.17$) and that are, except PW_RLE, comparable to what was labelled as ’mild’ or local damage in the beam numerical model;
  • 2.
    in the same locations, larger masses (M3, $\approx 900$ g) increase the MTMAC values up to $\approx 0.12-0.33$; again, except for a single case (in this instance, PW_CTE), these values are aligned with the ones caused by severe stiffness reduction in the beam model.
Summarising, it is clear that no absolute and universally-valid thresholds can be defined a priori, and in a complex system such as an entire aircraft, these depend on the specific location under scrutiny. Nevertheless, these values represent some insight about which orders of magnitude one should expect from this index.
Figure 16: BAE Systems Hawk T1A aircraft: Scaled COMAC for the port wing identified via iLF from cases (a) #1, (b) 2, (c) 3, (d) 4, (e) 5, and (f) 6.
@@
Cases #2 and 3 (Figures 16(a) and 16(b)): from the two images, it is clear that the areas most influenced by the added mass near the centre of the trailing edge (TE) are near the tip and closer to the root at $z\approx$ -2 m. Notably, as the mass is increased from M1 to M3, the deviation in the mode shapes becomes more dominant near the tip, while retaining some deviation at the same near-root area.
@@
Cases #6 and 7 (Figures 16(e) and 16(f)): the most prominent areas where a deviation in $\bm{\phi}_{1-3}$ are seen are the near-tip region and the same close to the root section in cases #2 and 3. Notably, the maximum deviation for both cases is seen near the damaged area. However, when the mass is increased from M2 to M3, the maximum deviation point shifts slightly towards the root (from $x\approx$ -4 m to -3.5 m).
@@ 4.2.2 Real, panel-off, damage
In Figure 17, the MTMAC values for the panel-off cases (#8-12) are shown. Notably, the values for the panels removed at PW2-3 and 5 have a similar magnitude, while for the PW1 and PW4 cases, the MTMAC values are doubled. These larger deviations in $\omega_{1-3}$ and $\bm{\phi}_{1-3}$ from case #1 show that the latter locations have a higher sensitivity to damage.
@@
Figure 18 shows the scaled COMAC for all of the panel-off cases. It is important to see that for all different locations, different deviations are identified with the COMAC; however, the maximum COMAC never exactly matches the damaged area. This is discussed later in the article.
@@
Figure 18: BAE Systems Hawk T1A aircraft: Scaled COMAC for the port wing identified via iLF from cases (a) #10, (b) 11, (c) 12, (d) 13, and (e) 14.
4.3 Discussion
@@
For all cases, added mass and panel-off, the MTMAC successfully detects a change in the system modal parameters $\omega_{1-3}$ and $\bm{\phi}_{1-3}$. In particular, from the two subsets (added mass and panel-off), it is possible to see which locations are more sensitive to changes in the structure. The highest MTMAC values are recorded for cases #8 and 11, which, respectively, represent panel-off cases at locations near the root (PW1) and at mid-span (PW4). The other three panel-off cases show a similar MTMAC value, hence showing a similar sensitivity. Notably, all panels are roughly placed between the quarter chord and mid-chord. On the other hand, the added mass locations are close to either the LE or TE. For them, the most sensitive location to damage is case #5, near the root leading edge. In contrast, the location least sensitive to damage is at TE at mid-span. Nevertheless, in all cases, it is possible to identify the increase in the added mass with an increase in the MTMAC, as shown in Figure 15. This means that MTMAC achieves unequivocality in the damage assessment, which is what is lacking in carrying out this assessment through modal parameters. This is confirmed by the early identified results in Figure 13(a) and Table 4.
@@
With respect to damage localisation, this is directly achieved only for cases #5-7. Nevertheless, each couple of added mass cases is somewhat coherent in their COMAC values distribution and, most importantly, it is unequivocally different from the others. The latter also holds for panel-off cases. These findings indicate that the COMAC can be effectively used for damage location in complex structures, particularly when paired with a high-fidelity finite element model of the system under examination or an empirical measure to correlate the COMAC value distribution with the predicted damage location. Unfortunately, neither was available for the wing under scrutiny at the time of writing; this aspect is not addressed in this work, as it lies beyond its scope, i.e., introducing the use of iLF for MTMAC-based SHM in MIMO tests.
Before moving to the conclusions, it is important to address some limitations of the newly-proposed method. As mentioned, the identification and damage assessment framework is limited to the linear (or linearised) response of the system under scrutiny; this, however, is conventional in current GVT applications. Regarding the interpretation of modal deformations, it must be said that while COMAC helps, the interpretation of complex mode shape deviations remains challenging, and location accuracy may be reduced in highly complex structures. This is, in fact, a recognised limitation of all mode-shape-based approaches. Finally, for field test validation, only one real aircraft dataset was used, with limited damage scenarios. Broader validation (other structures, other types of damage) would be needed for further generalisation, even if iLF has already proved effective in other Aerospace [52], Mechanical [34] and even Civil Engineering [70] applications. Furthermore, no environmental effects have been considered in this work as the data used in this work was collected in a controlled experimental environment.
However, iLF proved capable of robustly identifying the (linear) modal parameters of the system, which MTMAC successfully employed for structural assessment. In this regard, mode shapes can be extracted indifferently from displacement, velocities, or acceleration time series; in this work, acceleration time histories were considered only because they are much more common in standard practice.
@@ 5 Conclusions
In this work, a novel vibration-based damage detection and localisation method was applied, targeting the linearised dynamics (the regime in which most SHM and modal analysis methods are validated) of a mechanical system. In particular, the proposal has been verified on a numerically simulated hollow beam and a complex, real-life aeronautical case study of a trainer jet aircraft.
A two-step procedure has been introduced, including the identification of the modal parameters with the improved Loewner framework (iLF) and the detection and quantification of damage with the modified total modal assurance criterion (MTMAC). This is the first instance that the iLF has been used in structural health monitoring and that the MTMAC has been employed as a standalone damage assessment index.
First, the proposed approach has been validated on a numerical, fully controlled case study. In this benchmark, no nonlinear joints or contact effects were modelled; instead, the hollow beam was designed as a purely linear system to isolate the performance of the method.
Then, for experimental validation, damage detection and localisation have been innovatively carried out on the multi-input multi-output vibration dataset of the BAE Systems Hawk T1A aircraft from the Laboratory for Verification and Validation at the University of Sheffield. The proposed method has been shown to retain generality within the LTI assumption, as it applies equally to SIMO and MIMO testing, as shown by moving from the numerically simulated beam to the experimental dataset.
These two analyses resulted in the following findings:
@@
    The iLF has been numerically and experimentally proven to detect small damage-related changes in modal parameters from multi-input multi-output (MIMO) case studies - either numerically-simulated multiple degrees of freedom systems and complex experimental data (the Hawk T1A jet trainer);
@@
    Damage was successfully identified in 11 cases of the recently introduced Hawk T1A aircraft MIMO dataset, a first in the literature. Supporting the methodology feasibility for real-world SHM campaigns;
  • 4.
    The lack of partial and/or full airframe damage detection applications and methods is addressed.
Concerning future directions, this study focused solely on System Identification, aiming at data-driven damage assessment. However, the procedure presented here can also be paired with calibrated numerical models to compare the results with those obtained by FEA, potentially to identify regions of stress concentration (for damage localisation) and/or quantify damage (for severity assessment). Furthermore, while the framework successfully differentiated between healthy and damaged cases, no absolute, universally valid threshold has been set to define a ‘critical’ or irreversible structural state, because these will necessarily be case-specific. In practice, such thresholds could be determined statistically from repeated baseline measurements or derived from FEA/digital twin models coupled with certification criteria. Integrating these aspects will be an important direction for the practical deployment of the proposed method.
@@ Appendix A Euler-Bernoulli beam mass and stiffness matrices
Conceptualisation, G.D., and M.C.; methodology, G.D., and M.C.; software, G.D.; validation, G.D. and M.C.; formal analysis, G.D.; investigation, G.D. and M.C.; resources, G.D, M.C., B.C., A.M., and OE.BM.; data curation, G.D. and M.C.; writing—original draft preparation, G.D. and M.C.; writing—review and editing, G.D, M.C., A.M., B.C., and OE.BM.; visualisation, G.D., M.C, and A.M.; funding acquisition, G.D, M.C., B.C., A.M., and OE.BM..
@@
The first author is supported by the grants for research activity of young PhD holders, part of the Universidad Carlos III de Madrid (UC3M) Own Research Program (Ayudas para la Actividad Investigadora de los Jóvenes Doctores, del Programa Propio de Investigación de la UC3M). Furthermore, the authors from UC3M disclosed receipt of the following financial support for the research, authorship, and/or publication of this article: This work has been supported by the Madrid Government (Comunidad de Madrid– Spain) under the Multiannual Agreement with the UC3M (IA_aCTRl-CM-UC3M). The second and fourth authors are supported by the Centro Nazionale per la Mobilità Sostenibile (MOST – Sustainable Mobility Center), Spoke 7 (Cooperative Connected and Automated Mobility and Smart Infrastructures), Work Package 4 (Resilience of Networks, Structural Health Monitoring and Asset Management). The third author gladly acknowledges funding as a Beatriz Galindo Distinguished Senior Researcher by the Spanish Government.
@@ Acknowledgements
The experimental and numerical results (Data file 1) supporting this work are openly available from the Zenodo repository at [https://zenodo.org/records/13866467]. In addition, this study used third-party experimental data (Data file 2) made available at [orda.shef.ac.uk/articles/dataset/BAE_T1A_Hawk_Full_Structure_Modal_Test/24948549] under a licence that the authors do not have permission to share:
@@
    Data file 1. Data supporting: Multiple input tangential interpolation-driven damage detection of a jet trainer aircraft;
@@
    Data file 2. BAE T1A Hawk Full Structure Modal Test.